OpenStudy (anonymous):

Find the derivative of f(x)=|x^2-3x+2|

OpenStudy (anonymous):

\[f(x)=|x^2-3x+2|=\sqrt{(x^2-3x+2)^2}\]Then, differentiate it.

OpenStudy (anonymous):

The thing that I was confused about was the fact that it's absolute value...so I have to make it piecewised...How do I figure out the piecewise derivative?

OpenStudy (anonymous):

Nope, you don't have to.. Make use of the fact that \(|a| = \sqrt{a^2} \) and then use chain rule, you can do it.

OpenStudy (anonymous):

Ok, but the answer that the professor wants from me is a piecewise function.

OpenStudy (anonymous):

With all the domains. Sorry.

OpenStudy (anonymous):

Hmm.. Then... Do you know when the function will be positive and negative?

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

For example, when x =3, x^2-3x+2 = 9-9+2 = 2 => Positive Oh.. for the ''function'', I mean for x^2 - 3x +2

OpenStudy (anonymous):

ok so will the function be positive for x>1?

OpenStudy (anonymous):

and positive for x>2

OpenStudy (anonymous):

and negative for 1<x<2

OpenStudy (anonymous):

What about when x is negative?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

When x is negative, is f(x) positive or negative?

OpenStudy (anonymous):

Oh i figured it out! Thank you!

OpenStudy (anonymous):

wait, so since the function is positive when x>2 or x<1, that means that for those 2 parts, the derivative will be 2x-3 and since the function is negative from 1<x<2, the derivative will be -2x+3. Right?

OpenStudy (anonymous):

Yes.

OpenStudy (raden):

i think, @anneflys want to find derivative and domain of derivative, right ?

OpenStudy (anonymous):

I already knew how to find the domain though.

OpenStudy (raden):

how do u get domain of derivative, if derivative not ready :)

OpenStudy (anonymous):

I figured it out. It's all good. Thanks.

OpenStudy (raden):

what's derivative of f(x) did u get

OpenStudy (raden):

show to me if u have cleared it, @anneflys

OpenStudy (anonymous):

It's already up there. "wait, so since the function is positive when x>2 or x<1, that means that for those 2 parts, the derivative will be 2x-3 and since the function is negative from 1<x<2, the derivative will be -2x+3. Right?"

OpenStudy (anonymous):

domain for 2x-3 is x>2 or x<1 and domain for -2x+3 is 1<x<2...

OpenStudy (raden):

abs function always be a positive for x real numbers, so dont think that

OpenStudy (anonymous):

The range of abs function is positive, but the domain can be negative.

OpenStudy (raden):

dont u look in the sqrt have square sign (.......)^2, what's mean this

OpenStudy (anonymous):

When x =-100 \[\sqrt{a^2} = \sqrt{(-100)^2} = |100|\] -100 can be an input of the function..

OpenStudy (raden):

|100| = 100

OpenStudy (anonymous):

|-100| = 100

OpenStudy (raden):

that's i meant :)

OpenStudy (anonymous):

But then it's the range that must be positive, at least \(\ge\)0 But not the domain!

OpenStudy (anonymous):

Domain is the value of x (the input value), range is the output value..

OpenStudy (raden):

i know, so it means range of f(x) satisfied for domain all x's real in other word, domain = x real numbers

OpenStudy (anonymous):

But I think the domain she needs is the domain for the corresponding piecewise function..

OpenStudy (raden):

question from me for u @RolyPoly can u determine what's the value of "x", to get f(x) not real ??

OpenStudy (anonymous):

Do you mean for the domain, she needs to write it like this: t <1 , 1<t<2, t>2 , where \(t\in \mathbb{R} \)?

OpenStudy (raden):

i just want do contradiction, to prove it

OpenStudy (anonymous):

Hmm.. wait.. t <1 , 1<t<2, t>2 ^ That doesn't include 1 and 2 :O